# Security flaws from time-varying active encoding in high-speed measurement-device-independent quantum key distribution

### v1

@@ Abstract
Quantum key distribution (QKD) can transmit secret keys with, in principle, information-theoretic security. However, bandwidth limitations in practical equipment threaten the security of high-speed (GHz) QKD systems. We propose and characterize a new side channel which arises when using active encoding. As an illustrative example, we focus on electro-optic phase modulation for polarization encoding at 1 GHz. We show that this side channel may reduce the maximum secure transmission distance by over 50% in a decoy state measurement-device-independent QKD protocol.
@@
Introduction.— Quantum key distribution is a method which allows two parties to share a secret key with information-theoretic, quantum-safe security Xu et al. (2020). This is commonly done through encoding information into a train of quantum optical states. Unfortunately, practical implementations of QKD often fail to strictly uphold the security requirements which are posed by QKD theory for information-theoretic security. Imperfections come in the form of side channels which may leak information to an eavesdropper as well as state preparation flaws (SPFs) which are errors in the prepared quantum states Xu et al. (2020). As evidenced by various quantum hacking experiments, side channels present significant security risks to QKD protocols, making their characterization and mitigation vital areas of research Xu et al. (2020, 2010); Gottesman et al. (2004); Tang et al. (2013); Zhao et al. (2008); Lydersen et al. (2010); Brassard (2005); Jain et al. (2014). The development of the measurement-device-independent QKD (MDI QKD) protocol marked a significant milestone in QKD research by eliminating all side channels at the detector Lo et al. (2012). However, source side channels remain a major issue requiring further research Lo et al. (2012); Xu et al. (2020). Source side channels and imperfections could significantly lower the secret key rate in QKD Currás-Lorenzo et al. (2023); Tamaki et al. (2014) and therefore deserve our close attention.
The secret key rate per second is crucial for assessing QKD system performance. Rapid key transmission is practically useful, especially when using QKD alongside the one-time pad encryption technique, which requires the key length to match the message length for information-theoretic security Shannon (1949). As the secret key rate per pulse is fundamentally constrained by limits such as the PLOB bound Pirandola et al. (2017), enhancing the optical pulse transmission speed becomes pivotal for elevating the secret key rate. This has motivated a push towards high speed (GHz clock rate) QKD Roger et al. (2023); Boaron et al. (2018); Li et al. (2019); Grünenfelder et al. (2020); ichiro Yoshino et al. (2018); Woodward et al. (2021); Beutel et al. (2021); Chen et al. (2023); Lu et al. (2021). However, at high speeds, limited equipment bandwidth leads to the emergence of novel side channels at the source. The pattern effect is a notable example which has garnered significant interest ichiro Yoshino et al. (2018); Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018); Pereira et al. (2020).
Here, we propose and characterize a new side channel which may pose a significant threat to the security of high speed QKD systems - time-varying encoding. Consider the standard BB84 QKD protocol, which requires each prepared state to be randomly chosen from one of four specific qubit states which are required for information-theoretic security Xu et al. (2020). At high clock rates, using bandwidth-limited modulation equipment to encode optical states can result in encoding variations across the state duration/width. Therefore, the optical states possess a temporal degree of freedom tied to the encoding process, which serves as a side channel. We emphasize that this would be an issue in any discrete-variable QKD protocol using bandwidth-limited active modulation for encoding.
This has motivated our theoretical and experimental study of this newly identified side channel. As an illustrative example, we focus on active BB84 polarization encoding, which is often achieved using an electro-optic phase modulator. We present simulated and experimental time-varying polarization encoding profiles corresponding to a repetition rate of 1 GHz. Using a numerical security proof based on semidefinite programming Primaatmaja et al. (2019), we use our simulated and experimental results to obtain the asymptotic key rate for a decoy state MDI QKD protocol. Our results demonstrate a significant (over 50%) reduction in the maximum secure transmission distance due to the time-varying encoding side channel.
Quantum Optical Modelling.— Here, we derive an expression for the single photon component of a quantum optical coherent state with time-varying polarization encoding. This will be used to compute the single photon phase error rate (used to bound an eavesdropper’s information on the secret key in a decoy state protocol Lo et al. (2005)). We will use |$\ket{\psi^{i}_{x}}$ to represent this single photon component, where $i$ represents the basis choice and $x$ represents the bit choice.
@@
We start by expressing the time-varying polarization state of |$\ket{\psi^{i}_{x}}$. Note that active polarization encoding is typically achieved using an electro-optic phase modulator to modulate the phase along a specific polarization component. Without loss of generality, we consider an input (unmodulated) polarization state of |$\frac{\ket{H}+\ket{V}}{\sqrt{2}}$ (where |$\ket{H}$ and |$\ket{V}$ are orthogonal, linear polarization states). Now, let $\phi^{i}_{x}(t)$ refer to the time-varying phase change applied by the phase modulator to the |$\ket{V}$ component. The output time-varying polarization state would then be |$\frac{\ket{H}+e^{i\phi^{i}_{x}(t)}\ket{V}}{\sqrt{2}}$.
@@
Now, we consider the optical intensity profile of the quantum optical coherent state. We assume a Gaussian intensity profile. Hence, the temporal probability distribution of the single photon component would follow a Gaussian function, $f(t)=ae^{-\frac{t^{2}}{2c^{2}}}$. We choose $a$ such that $\int_{-\infty}^{\infty}f(t)=1$. The parameter $c$ is related to the full width at half maximum (FWHM) of the optical pulse (recall that the FWHM of a Gaussian pulse is $2\sqrt{2\ln 2}c$).
Hence, |$\ket{\psi^{i}_{x}}$ can be written as follows (leaving out the global phase term, which we assume in this work to be time-independent).
@@
Time-Varying Encoding is not Loss-Tolerant.— We can intuitively understand the level of information leakage inherent in |$\ket{\psi^{i}_{x}}$ using the formulation described in Currás-Lorenzo et al. (2023). We start by breaking down the state as follows.
|$\ket{\psi^{i}_{x}}=\sqrt{1-\epsilon^{i}_{x}}\ket{\theta^{i}_{x}}+\sqrt{\epsilon^{i}_{x}}\ket{\theta^{i\perp}_{x}}$ (2)
Here, |$\ket{\theta^{i}_{x}}$ is the closest (in terms of fidelity) qubit state to |$\ket{\psi^{i}_{x}}$. Here, we assume the form of such a state to match Equation 1 with one key difference - the polarization is not time-dependent (which results in separability between the temporal degree of freedom and polarization degree of freedom). Hence, we can express |$\ket{\theta^{i}_{x}}$ as follows, where $\theta^{i}_{x}$ represents a time-independent phase change applied by the phase modulator ($\theta^{i}_{x}$ is chosen such that $|\braket{\theta^{i}_{x}}{\psi^{i}_{x}}|^{2}$ is maximal).
@@
The remaining component, |$\ket{\theta^{i\perp}_{x}}$ effectively represents the side channel within the transmitted state and is orthogonal to |$\ket{\theta^{i}_{x}}$. The parameter $\epsilon^{i}_{x}$ represents the level of deviation of |$\ket{\psi^{i}_{x}}$ from a qubit state due to time-varying encoding. It can be expressed as follows.
@@
Note that the imperfection introduced by time-varying encoding differs from a simple SPF, which would only introduce a polarization error while preserving the encoded state as a qubit. As demonstrated by the loss-tolerant protocol for bounding secret key rates in QKD, SPFs do not render a QKD protocol vulnerable to channel loss Tamaki et al. (2014); Pereira et al. (2019). However, this is not the case for the side channel introduced by a non-qubit component of the encoded state Currás-Lorenzo et al. (2023). Eve could potentially exploit channel loss to extract information about the encoded states through this side channel (by performing an unambiguous state discrimination attack, for example), posing significant threats to the secret key rate and maximum secure transmission distance Currás-Lorenzo et al. (2023). The impact of the side channel on security can be understood through the parameter $\epsilon^{i}_{x}$. We will later use this parameter in order to further understand and contextualize our secret key rate results.
Characterization of Time-Varying Phase Modulation.— Here, we acquire $\phi^{i}_{x}(t)$ (which we will refer to as “temporal phase profiles” or simply “phase profiles”) in order to concretely understand the level of time-variance in high-speed bandwidth-limited electro-optic phase modulation when using our state-of-the-art lab equipment.
We start by reviewing electro-optic phase modulation Gee et al. (1983). An electro-optic phase modulator works through a traveling wave mechanism in which an optical signal and traveling RF signal propagate together through the modulator with their speeds matched. The resulting temporal phase profile of the optical signal is proportional to the temporal voltage profile of the RF signal. The voltage required to create a $\pi$ phase change is referred to as $V_{\pi}$ (more generally, we will use $V_{\phi}$ to refer to the voltage required to create a $\phi$ phase change).
Recall that in this paper, we consider BB84 encoding. Hence, we are interested in the temporal phase profiles resulting from the following phase modulations - $\pi/2$, $\pi$, and $-\pi/2$. The width of the phase modulation window (i.e. the width of the voltage pulse used to modulate phase) must be adequately larger than the width of the optical pulse being modulated. Here, we consider a 50 ps wide optical pulse and a 200 ps wide phase modulation window, typically used at GHz clock rates Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018); ichiro Yoshino et al. (2018); Li et al. (2019). Ideally, the voltage pulse used to modulate phase should be square (or have a flat modulation region at the peak), resulting in time-invariant modulation. However, the bandwidth limitations of the arbitrary waveform generator (AWG) and RF amplifier (which is often necessary to create voltage pulses with the required amplitude) result in distorted voltage pulses, deviating from a perfect square shape.
While limited bandwidth is a notable imperfection impacting the time-variance of phase modulation, other equipment-specific factors, such as nonlinearity and the frequency response function, may also contribute. This motivates us to provide optimistically simulated phase profiles in addition to our measured phase profiles as a generalized baseline. The simulated phase profiles solely consider the bandwidth limitations of the AWG, RF amplifier, and electro-optic modulator. We optimistically approximate the frequency response of each component using a simple Butterworth filter of order 2, with -3 dB points of 25 GHz, 11 GHz, and 15 GHz, respectively (corresponding to the equipment specifications). Butterworth filters are an optimistic choice, as their frequency response in the passband is relatively flat. Furthermore, at an order of 2, the frequency response falls off relatively gradually past the passband (in comparison to the frequency response of our state-of-the-art equipment) key (2017); drv (2023); jds (2004). These filters were created and applied to a 200 ps wide square pulse in order to generate the simulated profile (using the SciPy signal processing class) Contributors (2022). The result can be seen in Figure 1.
Figure 1: Simulated $\pi$ phase profile (the simulated ±$\pm\pi/2$ phase profiles are simply a scaled version of this). This was generated by applying Butterworth filters to a 200 ps wide square pulse. The Butterworth filters represent the AWG, amplifier, and electro-optic modulator. The optical pulse intensity profile assumed in this work (50 ps FWHM Gaussian pulse) is also shown in green.
We now move to our experimentally acquired temporal phase profiles. The experimental setup for generating and measuring the phase profiles is shown in Figure 2. The 25 GHz AWG was set to generate 200 ps wide (FWHM) square voltage pulses (the pulse repetition frequency was set to 10 MHz to reduce pulse correlations down to negligible levels). This pulse train was then amplified by an 11 GHz amplifier so that the voltage pulses have the necessary amplitude to generate the desired phase change. Using this setup, voltage pulses with an amplitude of $V_{\pi}$, $V_{\pi/2}$, and $V_{-\pi/2}$ were generated in order to acquire the $\pi$, $\pi/2$, and $-\pi/2$ temporal phase profiles. For our electro-optic modulator, $V_{\pi}$, $V_{\pi/2}$, and $V_{-\pi/2}$ are approximately 5.5 V, 2.75 V, and -2.75 V, respectively.
Figure 2: Experimental setup for generating and measuring phase profiles. The 25 GHz Keysight M8195A AWG (AWG) and 11 GHz ixBlue DR-VE-10-MO RF amplifier are used to generated voltage pulses which are sent to the JDSU OC-192 intensity modulator (IM). This intensity modulator acts as a phase modulator with a built in Mach-Zehnder interferometer to convert phase changes to intensity changes. The optical intensity profile (related to the optical phase profile through Equation 5) is collected using the Newport D-8ir 8 ps photodiode (PD) and the 12 GHz Agilent DSO81204A oscilloscope (OSC) after being amplified by a Thorlabs semiconductor optical amplifier (OA). The OA and PD were operated within their linear response region. Yellow indicates an optical fiber connection.
The $\pi$, $\pi/2$, and $-\pi/2$ temporal phase profiles were each generated by using the corresponding amplified voltage pulses to drive an electro-optic intensity modulator. An electro-optic intensity modulator uses electro-optic phase modulation in conjunction with Mach-Zehnder interferometry in order to modulate intensity. The phase change ($\phi$) applied by the intensity modulator is related to the normalized output intensity ($I$) through Equation 5 when the intensity modulator is biased at a minimum. This relationship can be used to determine a phase profile from its corresponding intensity profile.
$\phi=2cos^{-1}(\sqrt{1-I})$ (5)
Figure 3: a),b) Voltage pulses used for polarization/phase modulation as well as the corresponding phase profiles. Cubic spline interpolations of the profiles are shown using a solid line. The optical pulse intensity profile assumed in this work (50 ps FWHM Gaussian pulse) is also shown in green. Each profile was measured 10 times over an extended time period of several minutes (each time, oscilloscope averaging was used and set to 1024 to average out random noise, such as electronic noise) - the error bars represent the standard deviation in the 10 trials. c) Scaled version of a) demonstrating the difference between the voltage profiles used to modulate the two states. d) Scaled version of b) demonstrating the difference in temporal profiles between the two states.
The resulting temporal phase profiles are shown in Figure 3 (excluding the $-\pi$/2 profile, which we found to be nearly identical to the $\pi/2$ profile in shape). The corresponding voltage profiles are also shown. Evidently, the $\pi$ and $\pi/2$ phase profiles have a significant level of temporal variation (especially in comparison to the optimistically simulated phase profile from Figure 1, as expected). Interestingly, they do not have the same shape. We see the same effect in the corresponding voltage profiles. Likely, this results from the nonlinear (with respect to input voltage) response of the amplifier near its 1 dB compression point com of 5.6 V drv (2023).
In the next section, we will use our experimentally measured phase profiles in order to demonstrate the significant impact which time-varying polarization encoding has on the secret key rate in a decoy state MDI QKD protocol. We will also use our optimistically simulated phase profile to illustrate the extent of the impact, focusing solely on bandwidth limitations and excluding other practical considerations such as nonlinearity and imperfect frequency response.
Key Rate Results.— A symmetric polarization encoding (BB84) decoy state MDI QKD protocol with intensity settings of 0.05, 0.1, and 0.6 (near typical values Ma et al. (2012)) was assumed in the key rate analysis. Without loss of generality, we consider a projection onto the |$\ket{\Psi^{+}}$ Bell state as a passing measurement. A fiber loss of 0.2 dB/km and a dark count rate of $10^{-6}$/pulse was assumed. Moreover, the following simplifying assumptions were made to isolate specifically for the impact of time-varying encoding on the secret key rate. Although we neglect certain imperfections, they can be considered in a straightforward manner (unless stated otherwise) by deriving the observables used in the key rate formula accordingly and/or modifying the quantum optical model for the transmitted states.
  • Perfect polarization alignment to the electro-optic phase modulator (halfway between the fast and slow axes). Realistically, a certain degree of misalignment would be present.
  • A Gaussian optical pulse profile.
  • No temporal variation of global phase within the phase randomization window.
  • The temporal phase profiles of Bob’s states are the same as the temporal phase profiles of Alice’s states. In reality, Bob’s equipment specifications may be different from Alice’s equipment specifications. This may result in Bob’s states having different temporal phase profiles.
In our previous work Bourassa et al. (2022), we demonstrate the usage of a tight numerical key rate analysis technique based on semidefinite programming Primaatmaja et al. (2019) to deal with a specific time-dependent side channel. We show that the unique ability of the proof technique to deal with the exact time-dependent nature of the transmitted states (rather than a simplified, time-independent pessimistic approximation) results in a significantly tighter bound on the secret key rate. Hence, we chose to employ this same technique to analyze the impact of time-varying encoding on the secret key rate. Details regarding the technique can be found in the Supplementary Material.
At a high level, this technique uses the inner products between the transmitted states (i.e the Gram matrix of the transmitted states) to bound Eve’s information. This allows us to bound the secret key rate. Recall that here, the transmitted states can be expressed as per Equation 1, which we use to compute the inner products.
Using this technique, we simulate the secret key rate in four cases. The results can be seen in Figure 4.
  • Assuming the experimentally measured phase profiles shown in Figure 3b. Cubic spline interpolations (generated using the SciPy interpolation class) were used to represent the profiles, as shown in Figure 3b Contributors (2022). This allows the Gram matrix to be computed more reliably.
  • Assuming the optimistically simulated phase profile from Figure 1.
  • Assuming square phase profiles with a phase corresponding to the average phase of the experimentally measured phase profiles. We define the average phase ($\gamma_{avg}$) of an experimentally measured phase profiles ($\gamma(t)$) as follows.
@@
    Recall that $f(t)$ represents the Gaussian optical intensity profile of the transmitted states.
  • Assuming ideal square BB84 profiles.
Discussion.— In Figure 4, we see that the experimentally collected phase profiles lead to a significant (over 50%) reduction in the maximum secure transmission distance, relative to the ideal case. Since the simulated phase profiles are optimistic and do not consider any imperfections beyond bandwidth limitations, they perform slightly better. However, even in this relatively optimistic case, the maximum secure transmission distance is reduced by over 50%.
@@
Comparatively, despite having the same average encoding error as the experimentally collected phase profiles, the square profiles only cause a marginal reduction in the maximum secure transmission distance. Here comes a key conceptual point. A square phase profile implies that transmitted states satisfy the qubit assumption and hence do not carry a side channel. This results in a higher maximum secure transmission distance, as the protocol is now loss-tolerant Currás-Lorenzo et al. (2023).
In contrast, a temporally varying phase profile violates the qubit assumption. This provides additional dimensions to the Hilbert space and may enable an eavesdropper to perform new attack strategies. For this reason, the secure key is substantially reduced.
@@
Figure 4: Key rate results (with respect to the distance between Alice/Bob and the central relay) assuming an optical intensity profile with a 50 ps width (FWHM). In blue is the key rate plot which comes from using the experimentally collected phase profiles (from Figure 3). In orange is the key rate plot which comes from using the simulated phase profile (from Figure 1). In green is the key rate plot which comes from assuming square phase profiles with a time-independent error. This error corresponds to the average error of the experimental profiles. In red is the key rate plot which comes from assuming ideal (square) phase profiles.

### v2

# Hidden multi-dimensional modulation side channels in quantum protocols

@@ Abstract
Quantum protocols including quantum key distribution and blind quantum computing often require the preparation of quantum states of known dimensions. Here, we show that, rather surprisingly, hidden multi-dimensional modulation is often performed by practical devices. This violates the dimensional assumption in quantum protocols, thus creating side channels and security loopholes. Our work has important impacts on the security of quantum cryptographic protocols.
@@
Introduction.— Quantum protocols such as quantum key distribution (QKD) and blind quantum computing allow for information-theoretic, quantum-safe security Xu et al. (2020); Fitzsimons (2017). Thanks to measurement-device-independent QKD (MDI QKD), all side channels on the detection end can be completely eliminated Lo et al. (2012). Therefore, the source, as well as device-independent (DI) QKD are the last frontiers of investigation for security, with DI QKD remaining impractical due to its demanding requirements, such as near-perfect detector efficiencies Zapatero et al. (2023a); Xu et al. (2020); Nadlinger et al. (2022); Zhang et al. (2022). Practical implementations of non-DI protocols involve quantum state preparation that deviates from the theoretical requirements for security Xu et al. (2020). The loss-tolerant protocol illustrates that as long as the prepared quantum states satisfy a dimensional assumption, minor errors will not have a significant impact on security – otherwise, a side channel will be opened, causing a significant reduction to the maximum secure transmission distance Tamaki et al. (2014); Currás-Lorenzo et al. (2023).
Side channels and security loopholes in general are a point of significant interest in the literature for quantum cryptographic protocols, including QKD, quantum bit commitment, quantum coin flipping and position-based quantum cryptography Xu et al. (2020, 2010); Gottesman et al. (2004); Tang et al. (2013); Zhao et al. (2008); Lydersen et al. (2010); Brassard (2005); Jain et al. (2014). Notable examples include the pattern effect ichiro Yoshino et al. (2018); Xing et al. (2024), phase-remapping attack Fung et al. (2007); Xu et al. (2010), and the laser seeding attack Sun et al. (2015). In this Letter, we point out, rather surprisingly, that the commonly made dimensional assumption is often violated during the modulation process in quantum state preparation. We call the resulting class of side channels “hidden multi-dimensional modulation side channels.” Modulation is an important part of quantum state preparation (for example, during encoding). We show that in practical experiments, the modulation process may be tied to “hidden” degrees of freedom, causing a violation of the dimensional assumption. As an illustrative example, we focus on the case of time-varying polarization (BB84) encoding in MDI QKD. By time-varying encoding, we are referring to encoding which varies over the duration of the quantum optical state being encoded. Here, the “hidden” degree of freedom is time. However, our concept is general and applies to other degrees of freedom. While we focus on QKD for illustrative purposes, we emphasize that this issue would impact quantum cryptographic protocols in general.
The Three-State Protocol.— To understand the importance of the dimensional assumption in quantum protocols, let us start with the simple three-state QKD protocol Fung and Lo (2006) consisting of the states |$\ket{0}$, |$\ket{1}$, and |$\ket{+}={1\over\sqrt{2}}(\ket{0}+\ket{1})$. Notice that these states span two dimensions.
Practically, imperfections in preparing these states pose a security concern. Loss-tolerant QKD tells us that minor imperfections in the states do not affect the security of the protocol assuming that the states lie in a two-dimensional Hilbert space (i.e. the states satisfy the “qubit assumption”) Tamaki et al. (2014); Pereira et al. (2019). Such imperfections are referred to as “qubit flaws” Currás-Lorenzo et al. (2023). More specifically, the QKD protocol will remain robust to channel loss (“loss-tolerant”) Tamaki et al. (2014). However, we emphasize that this would not be the case if another dimension is present. We will see this point more clearly in the following protocol.
A Flawed Three-State Protocol.— Let us introduce a third dimension, |$\ket{2}$, such that |$\ket{0}$, |$\ket{1}$, and |$\ket{2}$ form an orthonormal basis. We now consider a flawed three-state protocol wherein the states lie in three dimensions, rather than two. As before, our conjugate basis consists of the state $\ket{+}$. However, our computational basis consists of |$\ket{0}$ and a flawed state, |$\ket{1^{\prime}}$. This flawed state takes the following form:
|$\ket{1^{\prime}}={\sqrt{1-\epsilon}}\ket{1}+{\sqrt{\epsilon}}\ket{2}.$ (1)
With some small amplitude, $\sqrt{\epsilon}$, |$\ket{1^{\prime}}$ lies in a third dimension, which is clearly a violation of the qubit assumption. Since the assumption is that Bob’s measurement is either inaccessible or oblivious to this third dimension, Alice’s inadvertent multi-dimensional modulation opens up a fatal attack strategy known as the unambiguous state discrimination (USD) attack Currás-Lorenzo et al. (2023); Ivanovic (1987). In this attack, Eve’s measurement either provides no information about the state or perfectly identifies it.
We now illustrate a potential USD attack for our flawed three-state protocol. Here, Eve performs a POVM with two elements.
$P_{1}=\ket{0}\bra{0}+\ket{1}\bra{1}$ | (2)
and
$P_{2}=\ket{2}\bra{2}.$ | . (3)
Notice that the outcome of the second element allows Eve to identify the state |$\ket{1^{\prime}}$ unambiguously. We can also construct POVM measurements to identify the other two states unambiguously. Furthermore, Eve can exploit practical channel loss to mask her measurement. This is why state preparation imperfections which violate the qubit assumption are not loss-tolerant. Hence, the USD attack is a significant threat to the security of quantum protocols such as QKD. While we used a three-state protocol to illustrate the USD attack, it applies to any number of states.
Hidden Multi-Dimensional Modulation Side Channel.— Now that we have reviewed how violating the qubit assumption can lead to the fatal USD attack, we can explain the hidden multi-dimensional modulation side channels which arise from this violation. We start by presenting a quantum optical model to elucidate the concept of these side channels, using our case example of time-varying polarization (BB84) encoding. Consider the single-photon component, |$\ket{\psi^{i}_{x}}$, of an encoded quantum state with time-varying polarization, where $i\in\{0,1\}$ represents the bit choice and $x\in\{0,1\}$ represents the basis choice.
@@
We first focus on the polarization degree of freedom. Note that active polarization encoding is typically achieved using an electro-optic phase modulator Gee et al. (1983) to modulate the phase along a specific polarization component. Without loss of generality, we consider an input (unmodulated) polarization state of |$\frac{\ket{H}+\ket{V}}{\sqrt{2}}$. Here, |$\ket{H}$ and |$\ket{V}$ are orthogonal and linear polarization states, with |$\ket{V}$ being the component along which phase modulation occurs. Hence, the polarization degree of freedom for a modulated state would take the following form:
|$\frac{\ket{H}+e^{i\phi^{i}_{x}}\ket{V}}{\sqrt{2}}$ (4)
As we consider BB84 encoding, $\phi^{i}_{x}$ would take on the following nominal values – 0, $\pi$, $\pi/2$, $-\pi/2$. However, in practice, there are two factors to take note of. Firstly, the quantum optical state being modulated has a finite duration. Secondly, due to the bandwidth limitations of the equipment driving the electro-optic modulator, the applied phase change would not be constant over the duration of the quantum optical state. Hence, the modulation process involves a “hidden” temporal degree of freedom.
@@
We now incorporate this temporal degree of freedom into our model. Let $f(t)$ represent the intensity profile of the quantum optical state. Here, we assume a Gaussian intensity profile. Hence, the temporal probability distribution of the single-photon component would follow a Gaussian function, $f(t)=ae^{-\frac{t^{2}}{2c^{2}}}$. We choose $a$ such that $\int_{-\infty}^{\infty}f(t)=1$. The parameter $c$ is related to the full width at half maximum (FWHM) of the optical pulse ($2\sqrt{2\ln 2}c$). Now, let $\phi^{i}_{x}(t)$ refer to the time-varying phase change applied by the phase modulator to the |$\ket{V}$ component. We can then represent |$\ket{\psi^{i}_{x}}$ as a summation of time-dependent polarization states over infinitesimal time-bins, |$\ket{t}$, as follows (leaving out the global phase term, which we assume in this work to be time-independent):
@@
It is clear that there is no separability between the time and polarization (encoding) degrees of freedom in (5). In other words, the temporal degree of freedom is coupled to the encoding degree of freedom. We emphasize that while our model specifically considers polarization and time, these can easily be replaced with other degrees of freedom. In general, |$\ket{H}$ and |$\ket{V}$ in (5) can be replaced by |$\ket{0}$ and |$\ket{1}$, where |$\ket{0}$ and |$\ket{1}$ can be spatial modes, binary time bins, frequency bins, etc. Similarly, |$\ket{t}$ in (5) can be a different degree of freedom. The hidden multi-dimensional modulation side channel persists as long as the encoding degree of freedom is coupled with another degree of freedom.
This additional degree of freedom increases the dimension of the Hilbert space spanned by the encoded states. As a result, the states may be linearly independent, opening up the possibility of an unambiguous state discrimination (USD) attack Dušek et al. (2000); Ivanovic (1987). By exploiting channel loss to discard inconclusive USD measurements without detection, an eavesdropper can identify every state that reaches the detector.
The model presented in (5) will be used to compute the secret key rate when the time-varying encoding side channel is present. More specifically, the model will be used to compute the single-photon phase error rate (used to bound an eavesdropper’s information on the secret key in a decoy state protocol Lo et al. (2005)).
We can intuitively understand the level of information leakage inherent in |$\ket{\psi^{i}_{x}}$ by breaking down the state as follows Currás-Lorenzo et al. (2023):
|$\ket{\psi^{i}_{x}}=\sqrt{1-\epsilon^{i}_{x}}\ket{\theta^{i}_{x}}+\sqrt{\epsilon^{i}_{x}}\ket{\theta^{i\perp}_{x}}.$ (6)
Note the similarity between (6) and (1). Here, |$\ket{\theta^{i}_{x}}$ is the closest (in terms of fidelity) qubit state to |$\ket{\psi^{i}_{x}}$. Here, we assume the form of |$\ket{\theta^{i}_{x}}$ to match (5) with one key difference – the polarization is not time-dependent (which results in separability between the temporal and polarization degrees of freedom). Hence, we can express |$\ket{\theta^{i}_{x}}$ as
@@
where $\theta^{i}_{x}$ represents a time-independent phase change applied by the phase modulator ($\theta^{i}_{x}$ is chosen such that $|\braket{\theta^{i}_{x}}{\psi^{i}_{x}}|^{2}$ is maximal). The remaining component, |$\ket{\theta^{i\perp}_{x}}$ is orthogonal to |$\ket{\theta^{i}_{x}}$ and causes a violation of the qubit assumption by increasing the dimension of the Hilbert space. (Note, the implicit assumption of (6) is that the qubit subspace in the larger Hilbert space is defined to be the polarization state with a constant temporal profile, i.e. |$\ket{\theta^{i}_{x}}$.) The parameter $\epsilon^{i}_{x}$ represents the level of deviation of |$\ket{\psi^{i}_{x}}$ from a qubit state due to time-varying encoding. It can be expressed as follows:
@@
The impact of the side channel on security can be understood through the parameter $\epsilon^{i}_{x}$. We will later use this parameter in order to further understand and contextualize our secret key rate results.
Impact of Time-Varying Encoding on the Secret Key Rate— Here, we concretely demonstrate the impact of a hidden multi-dimensional encoding side channel through key rate analysis, using our case example of time-varying polarization encoding. We employ a tight numerical key rate analysis technique based on semidefinite programming Primaatmaja et al. (2019) (see the Supplemental Material  [^1] See the Supplemental Material, which includes Refs. Ma et al. (2012); Bourassa et al. (2022); key (2017); drv (2023); jds (2004); Contributors (2022), for the key rate analysis technique and methods for acquiring practical temporal phase profiles. for further details). At a high level, this technique uses the Gram matrix of the transmitted states to bound Eve’s information. Recall that here, the transmitted states can be expressed as per (5), which we use to compute the inner products.
Using this technique, we simulate the secret key rate for two cases of the temporal phase profile, $\phi^{i}_{x}(t)$ – time-independent (square) and Gaussian (see Figure 1). The time-independent phase is set to be the average phase of the Gaussian phase profile. We define the average phase, $\gamma_{avg}$, of a phase profile, $\gamma(t)$ as follows:
@@
Recall that $f(t)$ represents the Gaussian optical intensity profile of the transmitted states. We also simulate the secret key rate assuming a time-independent phase with a perfect BB84 amplitude (ideal square). The secret key rate results can be found in Figure 2.
Figure 1: Normalized temporal phase profiles used in the key rate analysis (results shown in Figure 2). The dashed black line indicates a square (time-invariant) profile while the solid black line indicates a Gaussian (time-varying) profile. Both profiles have a FWHM of 400 ps. The optical intensity profile, $f(t)$ (assumed in the key rate analysis) is also shown in green. It is a Gaussian intensity profile with a FWHM of 50 ps.
@@
In Figure 2, we see that the Gaussian phase profile leads to a significant (approximately 50%) reduction in the maximum secure transmission distance, relative to the ideal case. Comparatively, despite having the same average encoding error as the Gaussian profile, the square profile only causes a marginal reduction in the maximum secure transmission distance, relative to the ideal case. Here comes a key conceptual point. A square phase profile implies that transmitted states satisfy the qubit assumption and hence do not carry the time-varying encoding side channel. This is assuming the optical intensity profile is negligible outside the modulation window, as is the case here. This results in a higher maximum secure transmission distance, as the protocol is now loss-tolerant Currás-Lorenzo et al. (2023). In contrast, a temporally varying phase profile violates the qubit assumption. This provides additional dimensions to the Hilbert space and may enable a USD attack. Hence, the secure key rate is substantially reduced.
@@
Figure 2: Key rate results (with respect to the distance between Alice/Bob and the central relay) for varying phase profiles. Blue: Gaussian phase profile (illustrated in Figure 1). Orange: Square (time-independent over the duration of the optical pulse) phase profiles. The amplitude is set to be the average phase of the Gaussian phase profile (see (9)). Green: Square phase profiles with ideal amplitudes.
Practical Temporal Phase Profiles — Here, we present key rate analysis results using practical (experimental and simulated) $\phi^{i}_{x}(t)$. This allows us to concretely understand the impact of time-varying encoding in realistic QKD systems. We focus on high-speed phase modulation, as QKD systems operating at GHz clock rates are most impacted by bandwidth limitations, which lead to time-varying modulation (recall that a square modulation signal requires infinite bandwidth). High-speed transmission enhances the practicality of QKD, sparking recent interest in this area Roger et al. (2023); Boaron et al. (2018); Li et al. (2019); Grünenfelder et al. (2020); ichiro Yoshino et al. (2018); Woodward et al. (2021); Beutel et al. (2021); Chen et al. (2023); Lu et al. (2021). Therefore, our study of this newly identified side channel is crucial.
We performed our experimental measurement of $\phi^{i}_{x}(t)$ using an electro-optic modulator of 15 GHz bandwidth, driven by an arbitrary waveform generator of 25 GHz bandwidth and amplifier of 11 GHz bandwidth. These bandwidths are typical of components used in today’s GHz QKD experiments ichiro Yoshino et al. (2018); Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018). While limited bandwidth is a notable imperfection impacting the time-variance of phase modulation, other equipment-specific factors such as nonlinearity and the frequency response function also contribute. Measurement artifacts may also affect the results. This motivates us to consider optimistically simulated phase profiles as a generalized baseline. This provides us with a practical lower bound on the time-variance of the profiles. We consider a 50 ps wide optical pulse and a 200 ps wide phase modulation window, typically used at GHz clock rates Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018); ichiro Yoshino et al. (2018); Li et al. (2019). To focus on the impact of time-variance breaking the qubit assumption, we mitigate the influence of correlations, a separate side channel, by collecting $\phi^{i}_{x}(t)$ at a low repetition rate of 10 MHz. Refer to the Supplemental Material for details regarding the simulation and experimental methods. Figures 1 and 3 in the Supplemental Material provide the resulting simulated and experimental temporal phase profiles, respectively.
Figure 3: Key rate results (with respect to the distance between Alice/Bob and the central relay) for different phase profiles assuming an optical intensity profile with a 50 ps FWHM. The phase profiles have a 200 ps FWHM. Blue: Experimentally collected phase profiles. Orange: Simulated phase profile. Green: Square (time-independent over the duration of the optical pulse) phase profiles. The amplitude is set to be the average phase of the experimental phase profile (see (9)). Red: Square phase profiles with ideal amplitudes.
In Figure 3, we see that the experimentally collected phase profiles lead to a significant (approximately 70%) reduction in the maximum secure transmission distance, relative to the ideal case. Since the simulated phase profiles are optimistic and do not consider any imperfections beyond bandwidth limitations, they perform slightly better. However, even in this relatively optimistic case, the maximum secure transmission distance is reduced by approximately 60%. Figure 3 also demonstrates the secret key rate when using square profiles with amplitudes corresponding to the average phase (see (9)) of the experimental profiles (square). As the square profile satisfies the qubit assumption, it does not cause a drastic reduction in the maximum secure transmission distance.